# Machine Learning-based Unfolding for Cross Section Measurements in the Presence of Nuisance Parameters

### v1

@@ 1 Introduction
Recently, a number of machine learning-based approaches have been proposed to address this problem (Arratia:2021otl; Huetsch:2024quz), and among those, one of the earliest methods proposed is OmniFold (Andreassen2020; Andreassen:2021zzk). OmniFold (OF) is a classifier-based algorithm that iteratively reweights simulated data to match the experimental data. At the population level, OmniFold is an Expectation-Maximization (EM) algorithm (10.2307/2984875), provided an infinite sample size and optimal classifier. OmniFold has been shown to be effective in high-dimensional settings and successfully applied to experimental data from the Large Hadron Collider (LHC) at CERN and other particle and nuclear physics experiments (H1:2021wkz; H1prelim-22-031; LHCb:2022rky; Komiske:2022vxg; H1:2023fzk; H1prelim-21-031; Song:2023sxb; Pani:2024mgy; CMS-PAS-SMP-23-008; ATLAS:2024xxl; ATLAS:2024jrp; Badea2025UnbinnedThrust; Canelli2025UnbinnedUnfolding).
@@
$p_{Y}(y)=\int_{\mathcal{X}}k_{\theta}(y,x)p_{X}(x)dx,$ (2)
where the response kernel depends on some nuisance parameters $\theta$. Without a correctly specified response kernel, the solution by OmniFold and other machine learning methods will be biased. Traditionally, this bias has been addressed by repeating the measurement using systematic variations of the detector response, which is an expensive and conservative step. In this paper, we approach this problem by proposing a new algorithm, called Profile OmniFold (POF), for unfolding in the presence of nuisance parameters. POF can be seen as an extension to the original OF algorithm, which iteratively reweights the simulated data, but at the same time simultaneously updates the nuisance parameters. This paper builds upon and expands the preliminary results presented in the NeurIPS Machine Learning and the Physical Sciences (ML4PS) workshop paper zhu2024multidimensional.
@@ 2 EM Algorithm for Unfolding
and $f^{(0)}(x)\geq 0$ for any $x\in\mathcal{X}$ and $\int_{\mathcal{X}}f^{(0)}(x)dx=1$. The solution after $k$ iteration is given by $f^{(k)}$. To the best of our knowledge, the earliest description of this algorithm appears in the work of kondor1983, which referred to it as the method of convergent weights. They did not derive the algorithm from the EM perspective, but instead based on the intuition that
@@
where $r^{(k)}(y):=\frac{p_{Y}(y)}{p_{Y}^{(k)}(y)}$ is the ratio of the smearing density and the updated density from the algorithm after $k$ iteration. Therefore, by constructing a sequence $r^{(k)}(y)$ that converges to one, the hope is that the corresponding $f^{(k)}$ will converge to $p_{X}$. Kondor presented the description of the algorithm along with a few examples, but did not establish the convergence property of the algorithm. Subsequently, multhi1987; multhi1989; multhi1992 connected the algorithm to the maximization of a concave functional, namely the population-level log-likelihood for a density function $f$ on $\mathcal{X}$
@@
or equivalently, minimization of the Kullback-Leibler divergence (multhi1989)
@@
with respect to $f$. In particular, under the assumption that the kernel $p(y|x)$ is strictly positive on the compact support $0,1]^{2}$, multhi1987 showed that if $f^{(k)}$ converges to some $\tilde{f}$ with respect to the $L^{1}$ norm, then $\tilde{f}$ is a maximizer of (6). Subsequently, multhi1992 showed that $p_{Y}^{(k)}$ converges uniformly to $p_{Y}$ if $p_{X}$ is strictly positive. Without assuming compact support, Chae2019 proved that $\text{KL}(p_{Y},p_{Y}^{(k)})\rightarrow\inf_{f}\text{KL}(p_{Y},\int p(y|x)f(x)dx)$, provided there exists a convergent sequence $(f_{*}^{(k)})_{k\geq 1}$ such that $\text{KL}(p_{Y},p_{Y}^{(k)})\rightarrow\text{KL}(p_{Y},\int p(y|x)f_{*}^{(k)}(x)dx)$. Beyond these results, eggemontLariccia1995; eggemontLariccia1997; eggemont1999 established similar convergence results for a smoothed version of the EM algorithm. More recently, Crucinio2023 analyzed the theoretical properties of the expectation maximization smoothing (EMS) scheme and proposed a particle algorithm as a sequential Monte Carlo method to approximate the EMS iteration.
Before these developments, part of the motivation for understanding the algorithm (3) comes from the well-known results by shepp1982maximum; doi:10.1080/01621459.1985.10477119, where they studied the discretized version of (1) as a model for image reconstruction in Positron Emission Tomography (PET). In particle physics, similar models have also been applied in binned unfolding, where measurements are binned into a histogram (or are naturally represented as discrete, e.g., in images) and the particle-level spectrum is likewise represented as a histogram.
@@ 2.1 Binned Unfolding
In the binned setting, the task is to estimate the true unknown histogram mean $\bm{\lambda}=[\lambda_{1},...,\lambda_{B}]$, where $\lambda_{j}=\int_{\mathcal{X}_{j}}p_{X}(x)dx$ for bins $\mathcal{X}_{1},...,\mathcal{X}_{B}$ at the particle level. The observed data are the detector-level histogram $\mathbf{n}^{*}=[n_{1}^{*},...,n_{D}^{*}]^{T}$, where $D$ is the number of bins at the detector level. Since events can be modeled as Poisson point processes, each bin count independently follows a Poisson distribution (Kuusela2012StatisticalII; Blobel:2203257).
@@
To obtain the maximum likelihood estimate, a classical approach has been the D’Agostini iteration (DAgostini:1994fjx). D’Agostini iteration can be viewed as an EM algorithm with early stopping (Kuusela2012StatisticalII), which is equivalent to the procedure originally proposed in shepp1982maximum. The same algorithm has also been known as the Richardson-Lucy algorithm (Richardson:72; 1974AJ.....79..745L). Specifically, starting from an initial guess $\bm{\lambda}^{(0)}>\bm{0}$, each component of $\hat{\bm{\lambda}}^{(k+1)}$ is updated iteratively by
@@
After $k$ iterations, the solution is given by $\hat{\bm{\lambda}}^{(k)}=(\hat{\lambda}_{1}^{(k)},...,\hat{\lambda}_{B}^{(k)})$. As $k\rightarrow\infty$, it can be shown that $\hat{\bm{\lambda}}^{(k)}$ converges to the maximum likelihood estimate of $\bm{\lambda}$ (doi:10.1080/01621459.1985.10477119). Also, since each step of the iteration increases the likelihood monotonically, stopping early in the iterations regularizes the solution. Moreover, comparing with the update rule in (9), the algorithm (3) can be viewed as a continuous analog of the discrete update in (9), assuming the full efficiency (i.e. $\sum_{i}K_{ij}=1$ for all $j$).
@@ 2.2 Unbinned Unfolding
Proposition 1 establishes that solving the optimization problem above yields the algorithm (3). This result (and similar variants) has been presented in multhi1987; Andreassen2020; Falcao2025HighDimensionalUnfolding. We provide the proof in supplementary material for completeness.
@@ 2.3 Machine-learning based Method for Unfolding: OmniFold
Although the EM algorithm (3) provides a principled approach to solving the inverse problem in (1), this is challenging to implement in practice for two key reasons: (1) the analytic forms of $p_{Y}$ and $p(y|x)$ are typically unknown in particle and nuclear physics experiments, and (2) both distributions may be high-dimensional, making them difficult to estimate. Recently, however, a line of machine learning-based unfolding methods has independently developed variants of algorithm (3) that circumvent these challenges and enable obtaining solutions to (1) even in high-dimensional settings. The first (and so far, only) one to be deployed to experimental data is OmniFold (Andreassen2020; Andreassen:2021zzk). The core idea of OmniFold is to use neural network classifiers to estimate the density ratios involved in the EM algorithm (3), thereby avoiding explicit estimation of $p_{Y}$ or $p(y|x)$. This is made possible by access to a set of Monte Carlo (MC) simulated data $\{X_{i}^{\prime},Y_{i}^{\prime}\}_{i=1}^{n}\sim q_{X,Y}$, which in particle and nuclear physics is routinely available from high-fidelity detector simulations that mimic the data-generating process. The key assumption here is that $q(y|x)=p(y|x)$, meaning the response kernel (or the forward operator) stays the same between the MC and experimental data. However, it should be noted that the marginal distributions $q_{X}$ and $p_{X}$ (and hence $q_{Y}$ and $p_{Y}$) are not assumed to be the same. For simplicity, we will omit the subscript $X$ and $Y$ in what follows if there is no confusion. Under this setting, OmniFold approaches the unfolding task as follows:
@@
Notice that combining these two steps yields the update (14). Moreover, since each step involves a density ratio, it can be estimated using a binary classifier without estimation of the marginal densities separately; see Section 2.4 for details. These two steps also have intuitive interpretations: the first step learns to reweight the detector-level MC density $\tilde{q}^{(k)}(y)$ in the current iteration to match the detector-level experimental density $p(y)$. The second step pulls back this density ratio to the particle level and updates the corresponding particle-level weights. In the next iteration, the updated weights are pushed forward to the detector level, and the process repeats. Here, pushing forward (or pulling back) refers to transferring the corresponding weights between paired particle-level and detector-level events. Further details are available in Andreassen2020.
@@ 2.4 Estimating density ratio through binary classification
Under a Bayes optimal classifier and as the sample size tends to infinity, $\hat{r}(x)\to r(x)$. Similar results can also be shown if the weights are associated with $x_{i}^{\prime}$ instead of $x_{i}$. For more details about density ratio estimation using classifiers, see, for example, Andreassen:2019nnm; Cranmer:2015bka; Chapter 4 of sugiyama2012density.
@@ 2.5 Other machine learning approaches
Although this work is primarily motivated by OmniFold, which is a classifier-based EM algorithm, other machine learning approaches have also been proposed for unbinned unfolding. Notably, another line of machine learning methods for unfolding uses generative models to learn the conditional distribution of the unfolded events given the observed data (Datta2019; Bellagente2020; Shmakov2023; Backes2024; Diefenbacher2024; Butter2025_DistributionMappingUnfolding; Butter2025_SPINUP; Butter2025_AnalysisReadyGenUnfold; Barman2025_SparticleML; Petitjean2025_GenerativeJets). In particular, it proceeds as follows: Initialize $p^{(0)}(x)=q(x),p^{(0)}(y)=q(y)$, then for iteration $k$,
@@
where $p^{(k)}(y)=\int r^{(k)}(x)q(x|y)q(y)dx$. Since $p^{(0)}(x)=q(x)$ and $p^{(0)}(y)=q(y)$, by induction we have $p^{(k)}(y)=\tilde{q}^{(k)}(y)$ for all $k\geq 0$. This shows that the update for $p^{(k)}(x)$ matches that of $\nu^{(k)}(x)$ in (14) with $p^{(k)}(x)=\nu^{(k)}(x)q(x)$. While the generative unfolding is still an EM algorithm in population level, most paper have been focusing on one step of training a generative model to learn $p(x|y)$ without iterating the procedure necessarily. Implementation-wise, various generative models have been explored in this context, including generative adversarial networks (Datta2019; Bellagente2020), diffusion models (Shmakov2023), normalizing flows (Backes2024), Schrödinger Bridges (Diefenbacher2024), and conditional flow matching (Petitjean2025_GenerativeJets).
@@ 3 Unfolding in the Presence of Nuisance Parameters in the Forward Operator
Related work includes the approach of Chan2023, which also performs unbinned unfolding with nuisance parameter profiling. Their method uses neural networks to directly maximize the Poisson log-likelihood function. While a significant step forward, this approach is limited to the case where the particle-level data are unbinned but the detector-level data are binned so that one can write down the explicit Poisson likelihood for bin counts. On the other hand, POF is completely unbinned at both the detector and particle levels, and does not assume any parametric model.
@@ 3.1 Algorithm
where $p_{0}(\theta)$ is a prior on $\theta$ to constrain the nuisance parameter, usually determined from auxiliary measurements. This is not strictly a Bayesian prior, but rather can be viewed as an optional likelihood term for the auxiliary measurements (cranmer2015practical). One example is the Gaussian likelihood, $\log p_{0}(\theta)=-\frac{(\theta-\bar{\theta})^{2}}{2\sigma_{0}^{2}}$.
@@
Figure 1: An overview of the POF algorithm. Portions of the image have been adapted from Andreassen2020 for the original OmniFold algorithm. In step 1, the current particle-level weights $\nu^{(k)}$ are pushed to the detector level with the current nuisance parameters $\theta^{(k)}$, which are used to compute the density ratio $r^{(k)}$. In step 2, the ratio $r^{(k)}$ is pulled pack to the particle level using the same nuisance parameters. In step 3, the nuisance parameters are updated based on the current weights $\nu^{(k)}$ and the density ratio $r^{(k)}$. The procedure is iterated for a fixed number of times.
@@
To address this issue, we propose to use a goodness-of-fit statistic based on the weighted accuracy of the step-1 classifier. Recall that in Step 1, a classifier is trained to estimate the density ratio $r^{(k)}(y)=\frac{p(y)}{\tilde{q}^{(k)}(y)}$. If the iterative procedure converges to the correct solution, the reweighted distribution $\tilde{q}^{(k)}(y)$ should match the observed distribution $p(y)$. Consequently, the classifier will be trying to discriminate two distributions that are nearly identical. In that case, the classifier’s accuracy should approach 0.5. Based on this observation, we propose the following goodness-of-fit statistic:

### v2

@@ 1 Introduction
Recently, a number of machine learning-based approaches have been proposed to address this problem (Arratia and others, 2022; Huetsch and others, 2024), and among those, one of the earliest methods proposed is OmniFold (Andreassen et al., 2020, 2021). OmniFold (OF) is a classifier-based algorithm that iteratively reweights simulated data to match the experimental data. At the population level, OmniFold is an Expectation-Maximization (EM) algorithm (Dempster et al., 1977), provided an infinite sample size and optimal classifier. OmniFold has been shown to be effective in high-dimensional settings and successfully applied to experimental data from the Large Hadron Collider (LHC) at CERN and other particle and nuclear physics experiments (H1 Collaboration, 2022b, a; Komiske et al., 2022; LHCb Collaboration, 2023; H1 Collaboration, 2023b, a; Song, 2023; Pani, 2024; CMS Collaboration, 2024; ATLAS Collaboration, 2024b, a; H1 Collaboration, 2024; ATLAS Collaboration, 2025; H1 Collaboration, 2025; Badea et al., 2025; Huang et al., 2025; Canelli et al., 2025).
@@
$p_{Y,\theta}(y)=\int_{\mathcal{X}}k_{\theta}(y,x)p_{X}(x)dx,$ (2)
where the response kernel depends on some nuisance parameters $\theta$. Without a correctly specified response kernel, the solution by OmniFold and other machine learning methods will be biased. Traditionally, this bias has been addressed by repeating the measurement using systematic variations of the detector response, which is an expensive and conservative step. In this paper, we approach this problem by proposing a new algorithm, called Profile OmniFold (POF), for unfolding in the presence of nuisance parameters. POF can be seen as an extension to the original OF algorithm, which iteratively reweights the simulated data, but at the same time simultaneously updates the nuisance parameters. This paper builds upon and expands the preliminary results presented in the NeurIPS Machine Learning and the Physical Sciences (ML4PS) workshop paper Zhu et al. (2024).
@@ 2 EM Algorithm for Unfolding
and $f^{(0)}(x)\geq 0$ for any $x\in\mathcal{X}$ and $\int_{\mathcal{X}}f^{(0)}(x)dx=1$. The solution after $k$ iteration is given by $f^{(k)}$. To the best of our knowledge, the earliest description of this algorithm appears in the work of Kondor (1983), which referred to it as the method of convergent weights. They did not derive the algorithm from the EM perspective, but instead based on the intuition that
@@
where $r^{(k)}(y):=\frac{p_{Y}(y)}{p_{Y}^{(k)}(y)}$ is the ratio of the smearing density and the updated density from the algorithm after $k$ iteration. Therefore, by constructing a sequence $r^{(k)}(y)$ that converges to one, the hope is that the corresponding $f^{(k)}$ will converge to $p_{X}$. Kondor presented the description of the algorithm along with a few examples, but did not establish the convergence property of the algorithm. Subsequently, Mülthei and Schorr (1987, 1989); Mülthei (1992) connected the algorithm to the maximization of a concave functional, namely the population-level log-likelihood for a density function $f$ on $\mathcal{X}$
@@
or equivalently, minimization of the Kullback-Leibler divergence (Mülthei and Schorr, 1989)
@@
with respect to $f$. In particular, under the assumption that the kernel $p(y|x)$ is strictly positive on the compact support $0,1]^{2}$, Mülthei and Schorr (1987, Theorem 8) showed that if $f^{(k)}$ converges to some $\tilde{f}$ with respect to the $L^{1}$ norm, then $\tilde{f}$ is a maximizer of (6). Subsequently, Mülthei (1992, Theorem 5) showed that $p_{Y}^{(k)}$ converges uniformly to $p_{Y}$ if $p_{X}$ is strictly positive. Without assuming compact support, Chae et al. (2019, Theorem 1) proved that $\text{KL}(p_{Y},p_{Y}^{(k)})\rightarrow\inf_{f}\text{KL}(p_{Y},\int p(y|x)f(x)dx)$, provided there exists a convergent sequence $(f_{*}^{(k)})_{k\geq 1}$ such that $\text{KL}(p_{Y},p_{Y}^{(k)})\rightarrow\text{KL}(p_{Y},\int p(y|x)f_{*}^{(k)}(x)dx)$. Beyond these results, Eggermont and Lariccia (1995, 1997); Eggermont (1999) established similar convergence results for a smoothed version of the EM algorithm. More recently, Crucinio et al. (2023) analyzed the theoretical properties of the expectation maximization smoothing (EMS) scheme and proposed a particle algorithm as a sequential Monte Carlo method to approximate the EMS iteration.
Before these developments, part of the motivation for understanding the algorithm (3) comes from the well-known results by Shepp and Vardi (1982); Vardi et al. (1985), where they studied the discretized version of (1) as a model for image reconstruction in Positron Emission Tomography (PET). In particle physics, similar models have also been applied in binned unfolding, where measurements are binned into a histogram (or are naturally represented as discrete, e.g., in images) and the particle-level spectrum is likewise represented as a histogram.
@@ 2.1 Binned Unfolding
In the binned setting, the task is to estimate the true unknown histogram mean $\bm{\lambda}=[\lambda_{1},...,\lambda_{B}]$, where $\lambda_{j}=\int_{\mathcal{X}_{j}}p_{X}(x)dx$ for bins $\mathcal{X}_{1},...,\mathcal{X}_{B}$ at the particle level. The observed data are the detector-level histogram $\mathbf{n}^{*}=[n_{1}^{*},...,n_{D}^{*}]^{T}$, where $D$ is the number of bins at the detector level. Since events can be modeled as Poisson point processes, each bin count independently follows a Poisson distribution (Kuusela, 2012; Blobel, 2011).
@@
To obtain the maximum likelihood estimate, a classical approach has been the D’Agostini iteration (D’Agostini, 1995). D’Agostini iteration can be viewed as an EM algorithm with early stopping (Kuusela, 2012), which is equivalent to the procedure originally proposed in Shepp and Vardi (1982). The same algorithm has also been known as the Richardson-Lucy algorithm (Richardson, 1972; Lucy, 1974). Specifically, starting from an initial guess $\bm{\lambda}^{(0)}>\bm{0}$, each component of $\hat{\bm{\lambda}}^{(k+1)}$ is updated iteratively by
@@
After $k$ iterations, the solution is given by $\hat{\bm{\lambda}}^{(k)}=(\hat{\lambda}_{1}^{(k)},...,\hat{\lambda}_{B}^{(k)})$. As $k\rightarrow\infty$, it can be shown that $\hat{\bm{\lambda}}^{(k)}$ converges to the maximum likelihood estimate of $\bm{\lambda}$ (Vardi et al., 1985). Also, since each step of the iteration increases the likelihood monotonically, stopping early in the iterations regularizes the solution. Moreover, comparing with the update rule in (9), the algorithm (3) can be viewed as a continuous analog of the discrete update in (9), assuming the full efficiency (i.e. $\sum_{i}K_{ij}=1$ for all $j$).
@@ 2.2 Unbinned Unfolding
Proposition 1 establishes that solving the optimization problem above yields the algorithm (3). This result (and similar variants) has been presented in Mülthei and Schorr (1987); Andreassen et al. (2020); Falcão and Takacs (2025). We provide the proof in supplementary material for completeness.
@@ 2.3 Machine-learning based Method for Unfolding: OmniFold
Although the EM algorithm (3) provides a principled approach to solving the inverse problem in (1), this is challenging to implement in practice for two key reasons: (1) the analytic forms of $p_{Y}$ and $p(y|x)$ are typically unknown in particle and nuclear physics experiments, and (2) both distributions may be high-dimensional, making them difficult to estimate. Recently, however, a line of machine learning-based unfolding methods has independently developed variants of algorithm (3) that circumvent these challenges and enable obtaining solutions to (1) even in high-dimensional settings. The first (and so far, only) one to be deployed to experimental data is OmniFold (Andreassen et al., 2020, 2021). The core idea of OmniFold is to use neural network classifiers to estimate the density ratios involved in the EM algorithm (3), thereby avoiding explicit estimation of $p_{Y}$ or $p(y|x)$. This is made possible by access to a set of Monte Carlo (MC) simulated data $\{X_{i}^{\prime},Y_{i}^{\prime}\}_{i=1}^{n}\sim q_{X,Y}$, which in particle and nuclear physics is routinely available from high-fidelity detector simulations that mimic the data-generating process. The key assumption here is that $q(y|x)=p(y|x)$, meaning the response kernel (or the forward operator) stays the same between the MC and experimental data. However, it should be noted that the marginal distributions $q_{X}$ and $p_{X}$ (and hence $q_{Y}$ and $p_{Y}$) are not assumed to be the same. For simplicity, we will omit the subscript $X$ and $Y$ in what follows if there is no confusion. Under this setting, OmniFold approaches the unfolding task as follows:
@@
Notice that combining these two steps yields the update (14). Moreover, since each step involves a density ratio, it can be estimated using a binary classifier without estimation of the marginal densities separately; see Section 2.4 for details. These two steps also have intuitive interpretations: the first step learns to reweight the detector-level MC density $\tilde{q}^{(k)}(y)$ in the current iteration to match the detector-level experimental density $p(y)$. The second step pulls back this density ratio to the particle level and updates the corresponding particle-level weights. In the next iteration, the updated weights are pushed forward to the detector level, and the process repeats. Here, pushing forward (or pulling back) refers to transferring the corresponding weights between paired particle-level and detector-level events. Further details are available in Andreassen et al. (2020).
@@ 2.4 Estimating density ratio through binary classification
Under a Bayes optimal classifier and as the sample size tends to infinity, $\hat{r}(x)\to r(x)$. Similar results can also be shown if the weights are associated with $x_{i}^{\prime}$ instead of $x_{i}$. For more details about density ratio estimation using classifiers, see, for example, Andreassen and Nachman (2019); Cranmer et al. (2015); Chapter 4 of Sugiyama et al. (2012).
@@ 2.5 Other machine learning approaches
Although this work is primarily motivated by OmniFold, which is a classifier-based EM algorithm, other machine learning approaches have also been proposed for unbinned unfolding. Notably, another line of machine learning methods for unfolding uses generative models to learn the conditional distribution of the unfolded events given the observed data (Datta et al., 2019; Bellagente et al., 2020; Shmakov et al., 2023; Backes et al., 2024; Diefenbacher et al., 2024; Butter et al., 2025a, b, c; Barman et al., 2025; Petitjean et al., 2025). In particular, it proceeds as follows: Initialize $p^{(0)}(x)=q(x),p^{(0)}(y)=q(y)$, then for iteration $k$,
@@
where $p^{(k)}(y)=\int r^{(k)}(x)q(x|y)q(y)dx$. Since $p^{(0)}(x)=q(x)$ and $p^{(0)}(y)=q(y)$, by induction we have $p^{(k)}(y)=\tilde{q}^{(k)}(y)$ for all $k\geq 0$. This shows that the update for $p^{(k)}(x)$ matches that of $\nu^{(k)}(x)$ in (14) with $p^{(k)}(x)=\nu^{(k)}(x)q(x)$. While the generative unfolding is still an EM algorithm in population level, most paper have been focusing on one step of training a generative model to learn $p(x|y)$ without iterating the procedure necessarily. Implementation-wise, various generative models have been explored in this context, including generative adversarial networks (Datta et al., 2019; Bellagente et al., 2020), diffusion models (Shmakov et al., 2023), normalizing flows (Backes et al., 2024), Schrödinger Bridges (Diefenbacher et al., 2024), and conditional flow matching (Petitjean et al., 2025).
@@ 3 Unfolding in the Presence of Nuisance Parameters in the Forward Operator
Related work includes the approach of Chan and Nachman (2023), which also performs unbinned unfolding with nuisance parameter profiling. Their method uses neural networks to directly maximize the Poisson log-likelihood function. While a significant step forward, this approach is limited to the case where the particle-level data are unbinned but the detector-level data are binned so that one can write down the explicit Poisson likelihood for bin counts. On the other hand, POF is completely unbinned at both the detector and particle levels, and does not assume any parametric model.
@@ 3.1 Algorithm
where $p_{0}(\theta)$ is a prior on $\theta$ to constrain the nuisance parameter, usually determined from auxiliary measurements. This is not strictly a Bayesian prior, but rather can be viewed as an optional likelihood term for the auxiliary measurements (Cranmer, 2015). One example is the Gaussian likelihood, $\log p_{0}(\theta)=-\frac{(\theta-\bar{\theta})^{2}}{2\sigma_{0}^{2}}$.
@@
Figure 1: An overview of the POF algorithm. Portions of the image have been adapted from Andreassen et al. (2020) for the original OmniFold algorithm. In step 1, the current particle-level weights $\nu^{(k)}$ are pushed to the detector level with the current nuisance parameters $\theta^{(k)}$, which are used to compute the density ratio $r^{(k)}$. In step 2, the ratio $r^{(k)}$ is pulled pack to the particle level using the same nuisance parameters. In step 3, the nuisance parameters are updated based on the current weights $\nu^{(k)}$ and the density ratio $r^{(k)}$. The procedure is iterated for a fixed number of times.
@@
To address this issue, we propose to use a goodness-of-fit statistic based on the weighted accuracy of the step-1 classifier. Recall that in Step 1, a classifier is trained to estimate the density ratio $r^{(k)}(y)=\frac{p(y)}{\tilde{q}^{(k)}(y)}$. If the iterative procedure converges to the correct solution, the reweighted distribution $\tilde{q}^{(k)}(y)$ should match the observed distribution $p(y)$. Consequently, the classifier will be trying to distinguish two distributions that are nearly identical. In that case, the classifier’s accuracy should approach 0.5. Based on this observation, we propose the following goodness-of-fit statistic: